OpenStudy (anonymous):

Half life of 1000 years. How long would it take to decay to 40% of that amount.

OpenStudy (anonymous):

too damn long

OpenStudy (anonymous):

solve \[.4=(\frac{1}{2})^{\frac{t}{1000}}\] for t

OpenStudy (anonymous):

via \[\ln(.4)=\frac{t}{1000} \ln(.5)\] \[t=1000\times \frac{\ln(.4)}{\ln(.5)}\]

OpenStudy (anonymous):

right mathgirl?

OpenStudy (anonymous):

totally

OpenStudy (anonymous):

hmmm. what would be the next step to get it in years

OpenStudy (anonymous):

whip out a calculator and type it in

OpenStudy (anonymous):

i did. its coming up undefined. let me try again

OpenStudy (anonymous):

i get 1322 rounded

OpenStudy (anonymous):

oh yes. i put it in wrong. ok. thanks.

OpenStudy (anonymous):

yw